The Department of Education would like to test the hypothesis that the average debt load of graduating students with a Bachelor's degree is equal to $17,000. A random sample of 34 students had an average debt load of $18,200. It is believed that the population standard deviation for student debt load is $4,200. The Department of Education would like to set α = 0.05. The critical sample means for this hypothesis test would be

Question options:

$14,118.9, $19,881.2

$14,839.1, $19,160.9

$15,588.2, $18,411.8

$16,279.7, $17,720.3
Answer #1

Critical value =\pm z_{\alpha/2}=\pm z_{0.025}=\pm 1.96

z=\frac{x-\mu}{\sigma/\sqrt{n}}

x=\mu+z*\sigma/\sqrt{n}

µ = 17000

\sigma=4200

x_L=17000-1.96*4200/\sqrt{34}=15588.2

x_U=17000+1.96*4200/\sqrt{34}=18411.8

Option-C) $15,588.2, $18,411.8